# Multimodal Variational Autoencoder for Low-cost Cardiac Hemodynamics Instability Detection

### v2

@@ 2 Methods
CXR Encoder: Our CXR encoder incorporates a convolutional neural network (CNN) consisting of three Convolutional (CONV) layers. It processes a CXR image $\mathbf{X}_{\text{CXR}}\in\mathbb{R}^{H\times W\times C}$, where $H$, $W$, and $C$ represent the height, width, and number of channels, respectively. Each CONV layer is configured with a $3\times 3$ kernel, a stride of $2$, and padding of $1$. We use channel depths of $16$, $32$, and $64$ in CONV layers. Following this, the data is flattened and passed through two Fully Connected (FC) layers to produce the mean ($\boldsymbol{\mu}_{\mathbf{{\text{CXR}}}}$) and variance ($\boldsymbol{\sigma^{2}}_{\mathbf{{\text{CXR}}}}$) vectors, which are crucial for defining the CXR image’s latent space representation and establishing its approximate posterior distribution parameters $q_{\phi_{\text{CXR}}}(\mathbf{z}_{\text{CXR}}|\mathbf{X}_{\text{CXR}})$.
@@
ECG Encoder: The ECG encoder comprises three one-dimensional (1D) CONV layers. It processes the sequential ECG signal $\mathbf{x}_{\text{ECG}}\in\mathbb{R}^{L}$, where $L$ denotes the length of the signal. Each 1D CONV layer is configured with a kernel size of $1\times 3$, a stride of $2$, and padding of $1$. We use channel depths of $16$, $32$, and $64$ in three CONV layers. Following this, the signal is flattened and processed through two FC layers, yielding the mean ($\boldsymbol{\mu}_{\mathbf{{\text{ECG}}}}$) and variance ($\boldsymbol{\sigma^{2}}_{\mathbf{{\text{ECG}}}}$) which are crucial for the ECG signal’s latent space representation and its approximate posterior distribution $q_{\phi_{\text{ECG}}}(\mathbf{z}_{\text{ECG}}|\mathbf{x}_{\text{ECG}})$.
@@
Multimodal Integration: In multimodal integration, we employ a Product of Experts (PoE) [7] approach to combine the approximate posterior distributions from the CXR and ECG encoders with a standard Gaussian prior $p(z)=\mathcal{N}(0,\mathbf{I})$ into a unified latent variable, effectively synthesizing the individual expert opinions, where $\mathbf{I}$ is the identity matrix. The combined mean $\boldsymbol{\mu}$ and variance $\boldsymbol{\sigma^{2}}$ of the latent variable [10] are computed as: $\boldsymbol{\mu}=(\sum_{m=1}^{M}\boldsymbol{\mu_{m}}/\boldsymbol{\sigma_{m}^{2}})/(\sum_{m=1}^{M}1/\boldsymbol{\sigma_{m}^{2}})$ and $\boldsymbol{\sigma^{2}}=1/(\sum_{m=1}^{M}1/\boldsymbol{\sigma_{m}^{2}})$, where $m$ represents the modality, ranging from 1 to $M$. This leads to the calculation of the latent space variable $\mathbf{z}$ using the reparameterization trick $\mathbf{z}=\boldsymbol{\mu}+\boldsymbol{\sigma}\cdot\boldsymbol{\epsilon}$, where $\boldsymbol{\epsilon}$ is drawn from $\mathcal{N}(0,\mathbf{I})$. This step allows the decoders to generate diverse yet consistent reconstructions, effectively establishing the approximate joint posterior distribution $q_{\phi_{\text{CXR, ECG}}}(\mathbf{z}_{\text{CXR, ECG}}|\mathbf{X_{\text{CXR}}},\mathbf{x_{\text{ECG}}})$ and reflecting a deep understanding of the merged input data.
@@
Decoder Design: The CXR and ECG decoders aim to reconstruct inputs from the unified latent space $\mathbf{Z}$, which contains both modality-specific latent variables $\mathbf{z}_{\text{CXR}}$, $\mathbf{z}_{\text{ECG}}$, and shared latent variable $\mathbf{z}_{\text{CXR,ECG}}$. For CXR image reconstruction, the CXR decoder $p_{\theta_{\text{CXR}}}(\mathbf{X}_{\text{CXR}}|\mathbf{z}_{\text{CXR}})$ is a CNN consisting of three transposed convolutional layers, effectively restoring the sample $\mathbf{X}_{\text{CXR}}$ from the latent variable $\mathbf{z}_{\text{CXR}}$. Similarly, the ECG decoder $p_{\theta_{\text{ECG}}}(\mathbf{x}_{\text{ECG}}|\mathbf{z}_{\text{ECG}})$ utilizes three 1D transposed convolutional layers to convert $\mathbf{z}_{\text{ECG}}$ into the precise temporal waveform $\mathbf{x}_{\text{ECG}}$. This reconstruction process leverages the inherent flexibility of the Gaussian distribution within the latent space, enabling the production of diverse, high-fidelity reconstructions by sampling various $\mathbf{z}$ points. For the unimodal streams, the decoder uses their respective modality-specific latent variables, but for the multimodal stream, the decoder uses both modality-specific and shared latent variables for reconstruction.
@@
Tri-stream Pre-training Strategy: For pre-training, we extend [12, 30] to introduce a tri-stream strategy that processes CXR and ECG data both separately and jointly to capture the modality-specific and shared features of each modality in the latent space. We pass three data streams through our $\text{CardioVAE}_{\text{X, G}}$ model: 1) CXR only, 2) ECG only, and 3) paired CXR and ECG, and calculate three separate losses for each stream by incorporating Evidence Lower Bound (ELBO) [14]. We define $\mathcal{L}_{\text{CXR}}$ for CXR, $\mathcal{L}_{\text{ECG}}$ for ECG, and $\mathcal{L}_{\text{(CXR,ECG)}}$ for joint loss, as follows:
$\mathcal{L}_{\text{CXR}}=\mathbb{E}_{q_{\phi_{\text{CXR}}}(\mathbf{z}_{\text{CXR}}|\mathbf{X}_{\text{CXR}})}[\lambda_{\text{CXR}}\log p_{\theta_{\text{CXR}}}(\mathbf{X}_{\text{CXR}}|\mathbf{z}_{\text{CXR}})]$
$-\beta D_{\text{KL}}[q_{\phi_{\text{CXR}}}(\mathbf{z}_{\text{CXR}}|\mathbf{X}_{\text{CXR}})||p(z)],$ (1)
$\mathcal{L}_{\text{ECG}}=\mathbb{E}_{q_{\phi_{\text{ECG}}}(\mathbf{z}_{\text{ECG}}|\mathbf{x}_{\text{ECG}})}[\lambda_{\text{ECG}}\log p_{\theta_{\text{ECG}}}(\mathbf{x}_{\text{ECG}}|\mathbf{z}_{\text{ECG}})]$
$-\beta D_{\text{KL}}[q_{\phi_{\text{ECG}}}(\mathbf{z}_{\text{ECG}}|\mathbf{x}_{\text{ECG}})||p(z)],$ (2)
$\mathcal{L}_{\text{(CXR,ECG)}}=\mathbb{E}_{q_{\phi_{\text{CXR}}}(\mathbf{z}_{\text{CXR}}|\mathbf{X}_{\text{CXR}})}[\lambda_{\text{CXR}}\log p_{\theta_{\text{CXR}}}(\mathbf{X}_{\text{CXR}}|\mathbf{z}_{\text{CXR}})]$
$+\mathbb{E}_{q_{\phi_{\text{ECG}}}(\mathbf{z}_{\text{ECG}}|\mathbf{x}_{\text{ECG}})}[\lambda_{\text{ECG}}\log p_{\theta_{\text{ECG}}}(\mathbf{x}_{\text{ECG}}|\mathbf{z}_{\text{ECG}})]$
$-\beta D_{\text{KL}}[q_{\phi_{\text{(CXR,ECG)}}}(\mathbf{z}_{\text{(CXR, ECG)}}|\mathbf{X_{\text{CXR}}},\mathbf{x_{\text{ECG}}})||p(z)],$ (3)
@@
where the first part $\mathbb{E}_{q_{\phi}}$ of the losses is the expected conditional log-likelihood of the data given the latent variable $\mathbf{z}$, indicating how well the model reconstructs the data. This part takes the reconstructed output from the decoder and finds the reconstruction error. The second part of the losses, $D_{\text{KL}}$, is the Kullback-Leibler (KL) divergence [2] between the approximate posterior (CXR or ECG or CXR-ECG) and the prior distribution $p(z)$, following the Gaussian distribution, over the latent variables, calculated in encoders and multimodal integration phase. These losses aim to maximize the likelihood of the data while minimizing the difference between the approximate posterior and the prior. Moreover, in our losses, we incorporate two important hyperparameters $\lambda$ [17] to balance the weights of modalities and $\beta$ [6] to balance the trade-off between reconstruction loss and KL divergence. In practice, $\lambda=1$ and $\beta$ are slowly annealed from 00 to $1$ to form a valid lower bound on the evidence [1]. After calculating three losses in Eq. (1-3) for the three streams, we combine them to obtain a total loss $\mathcal{L}_{\text{total}}$ as follows:
@@ 3 Experimental Results and Analysis
For fine-tuning, we froze the layers in the encoders and fine-tuned FC layers on the ASPIRE registry dataset. We used $128$ nodes in FC layer and used a dropout of $0.5$. We evaluated the prediction performance using $10$-fold cross-validation with a training and validation ratio of $80:20$. We set the learning rate to $0.001$ and the batch size to $32$, and trained the model for $50$ epochs. We used Area Under Receiver Operating Curve (AUROC) and accuracy metrics to assess classification performance. Moreover, we calculated the $p$-values of our best-performing model against other models to show the statistical significance of the results. For a fair comparison, we used the same training settings and data partitioning for comparing methods [16, 18, 23, 30]. An Nvidia RTX4090 GPU was used for all experiments. The implementation of all the models was carried out in Python (version $3.10$) with PyTorch [21]. The source code can be found at: https://github.com/Shef-AIRE/AI4Cardiothoracic-CardioVAE.
@@ 4 Conclusion and Future Work
4.0.1 Disclosure of Interests:
The authors have no competing interests to declare.

### v3

@@ 2 Methods
CXR Encoder: Our CXR encoder incorporates a convolutional neural network (CNN) consisting of three Convolutional (CONV) layers. It processes a CXR image $\mathbf{X}_{\text{CXR}}\in\mathbb{R}^{H\times W\times C}$, where $H$, $W$, and $C$ represent the height, width, and number of channels, respectively. Each CONV layer is configured with a $3\times 3$ kernel, a stride of $2$, and padding of $1$. We use channel depths of $16$, $32$, and $64$ in CONV layers. Following this, the data is flattened and passed through two Fully Connected (FC) layers to produce the mean ($\boldsymbol{\mu}_{\mathbf{{\text{CXR}}}}$) and variance ($\boldsymbol{\sigma^{2}}_{\mathbf{{\text{CXR}}}}$) vectors, which are crucial for defining the CXR image’s latent space representation and establishing its approximate posterior distribution parameters $q_{\phi_{\text{CXR}}}(z|\mathbf{X}_{\text{CXR}})$.
@@
ECG Encoder: The ECG encoder comprises three one-dimensional (1D) CONV layers. It processes the sequential ECG signal $\mathbf{x}_{\text{ECG}}\in\mathbb{R}^{L}$, where $L$ denotes the length of the signal. Each 1D CONV layer is configured with a kernel size of $1\times 3$, a stride of $2$, and padding of $1$. We use channel depths of $16$, $32$, and $64$ in three CONV layers. Following this, the signal is flattened and processed through two FC layers, yielding the mean ($\boldsymbol{\mu}_{\mathbf{{\text{ECG}}}}$) and variance ($\boldsymbol{\sigma^{2}}_{\mathbf{{\text{ECG}}}}$) which are crucial for the ECG signal’s latent space representation and its approximate posterior distribution $q_{\phi_{\text{ECG}}}(z|\mathbf{x}_{\text{ECG}})$.
@@
Multimodal Integration: In multimodal integration, we employ a Product of Experts (PoE) [7] approach to combine the approximate posterior distributions from the CXR and ECG encoders with a standard Gaussian prior $p(z)=\mathcal{N}(0,\mathbf{I})$ into a unified latent space, effectively synthesizing the individual expert opinions, where $\mathbf{I}$ is the identity matrix. The combined mean $\boldsymbol{\mu}$ and variance $\boldsymbol{\sigma^{2}}$ of the latent space [10] are computed as: $\boldsymbol{\mu}=(\sum_{m=1}^{M}\boldsymbol{\mu_{m}}/\boldsymbol{\sigma_{m}^{2}})/(\sum_{m=1}^{M}1/\boldsymbol{\sigma_{m}^{2}})$ and $\boldsymbol{\sigma^{2}}=1/(\sum_{m=1}^{M}1/\boldsymbol{\sigma_{m}^{2}})$, where $m$ represents the modality, ranging from 1 to $M$. This leads to the calculation of the latent space variable $\mathbf{z}$ using the reparameterization trick $\mathbf{z}=\boldsymbol{\mu}+\boldsymbol{\sigma}\cdot\boldsymbol{\epsilon}$, where $\boldsymbol{\epsilon}$ is drawn from $\mathcal{N}(0,\mathbf{I})$. This step allows the decoders to generate diverse yet consistent reconstructions, effectively establishing the approximate joint posterior distribution $q_{\phi_{\text{CXR,ECG}}}(z|\mathbf{X_{\text{CXR}}},\mathbf{x_{\text{ECG}}})$ and reflecting a deep understanding of the merged input data.
@@
Decoder Design: The CXR and ECG decoders aim to reconstruct inputs from the latent variable $\mathbf{z}$, sampled from the unified latent space learned by the PoE [7] approach. For CXR image reconstruction, the CXR decoder $p_{\theta_{\text{CXR}}}(\mathbf{X_{\text{CXR}}}|z)$ is a CNN consisting of three transposed convolutional layers, effectively restoring sample $\mathbf{X_{\text{CXR}}}$ from latent variable $z$. Similarly, the ECG decoder $p_{\theta_{\text{ECG}}}(\mathbf{x_{\text{ECG}}}|z)$ utilizes three 1D transposed convolution layer to convert $\mathbf{z}$ into precise temporal waveform $\mathbf{x_{\text{ECG}}}$. This reconstruction process leverages the inherent flexibility of the Gaussian distribution within the latent space, enabling the production of diverse, high-fidelity reconstructions by sampling various $\mathbf{z}$ points.
@@
Tri-stream Pre-training Strategy: For pre-training, we extend [30, 12] to introduce a tri-stream strategy that processes CXR and ECG data both separately and jointly to capture the unique and shared features of each modality in the latent space. We pass three data streams through our $\text{CardioVAE}_{\text{X,G}}$ model: 1) CXR only, 2) ECG only, and 3) paired CXR and ECG, and calculate three separate losses for each stream by incorporating Evidence Lower Bound (ELBO) [14]. We define $\mathcal{L}_{\text{CXR}}$ for CXR, $\mathcal{L}_{\text{ECG}}$ for ECG, and $\mathcal{L}_{\text{(CXR,ECG)}}$ for joint loss, as follows:
$\mathcal{L}_{\text{CXR}}=\mathbb{E}_{q_{\phi_{\text{CXR}}}(z|\mathbf{X}_{\text{CXR}})}[\lambda_{\text{CXR}}\log p_{\theta_{\text{CXR}}}(\mathbf{X}_{\text{CXR}}|z)]-\beta D_{\text{KL}}[q_{\phi_{\text{CXR}}}(z|\mathbf{X}_{\text{CXR}})||p(z)],$ (1)
$\mathcal{L}_{\text{ECG}}=\mathbb{E}_{q_{\phi_{\text{ECG}}}(z|\mathbf{x}_{\text{ECG}})}[\lambda_{\text{ECG}}\log p_{\theta_{\text{ECG}}}(\mathbf{x}_{\text{ECG}}|z)]-\beta D_{\text{KL}}[q_{\phi_{\text{ECG}}}(z|\mathbf{x}_{\text{ECG}})||p(z)],$ (2)
$\mathcal{L}_{\text{(CXR,ECG)}}=\mathbb{E}_{q_{\phi_{\text{CXR}}}(z|\mathbf{X}_{\text{CXR}})}[\lambda_{\text{CXR}}\log p_{\theta_{\text{CXR}}}(\mathbf{X}_{\text{CXR}}|z)]$
$+\mathbb{E}_{q_{\phi_{\text{ECG}}}(z|\mathbf{x}_{\text{ECG}})}[\lambda_{\text{ECG}}\log p_{\theta_{\text{ECG}}}(\mathbf{x}_{\text{ECG}}|z)]$
$-\beta D_{\text{KL}}[q_{\phi_{\text{(CXR,ECG)}}}(z|\mathbf{X_{\text{CXR}}},\mathbf{x_{\text{ECG}}})||p(z)],$ (3)
@@
where the first part $\mathbb{E}_{q_{\phi}}$ of the losses is the expected conditional log-likelihood of the data given the latent variable $z$, indicating how well the model reconstructs the data. This part takes the reconstructed output from the decoder and finds the reconstruction error. The second part of the losses, $D_{\text{KL}}$, is the Kullback-Leibler (KL) divergence [2] between the approximate posterior (CXR or ECG or CXR-ECG) and the prior distribution $p(z)$, following the Gaussian distribution, over the latent variables, calculated in encoders and multimodal integration phase. These losses aim to maximize the likelihood of the data while minimizing the difference between the approximate posterior and the prior. Moreover, in our losses, we incorporate two important hyperparameters $\lambda$ [17] to balance the weights of modalities and $\beta$ [6] to balance the trade-off between reconstruction loss and KL divergence. In practice, $\lambda=1$ and $\beta$ are slowly annealed from 00 to $1$ to form a valid lower bound on the evidence [1]. After calculating three losses in Eq. (1-3) for the three streams, we combine them to obtain a total loss $\mathcal{L}_{\text{total}}$ as follows:
@@ 3 Experimental Results and Analysis
For fine-tuning, we froze the layers in the encoders and fine-tuned FC layers on the ASPIRE registry dataset. We used $128$ nodes in FC layer and used a dropout of $0.5$. We evaluated the prediction performance using $10$-fold cross-validation with a training and validation ratio of $80:20$. We set the learning rate to $0.001$ and the batch size to $32$, and trained the model for $50$ epochs. We used Area Under Receiver Operating Curve (AUROC) and accuracy metrics to assess classification performance. Moreover, we calculated the $p$-values of our best-performing model against other models to show the statistical significance of the results. For a fair comparison, we used the same training settings and data partitioning for comparing methods [16, 23, 18, 30]. An Nvidia RTX4090 GPU was used for all experiments. The implementation of all the models was carried out in Python (version $3.10$) with PyTorch [21].